# StrucText-Eval: An Autogenerated Benchmark for Evaluating Large Language Model’s Ability in Structure-Rich Text Understanding

### v2

@@ 5 Analysis
CoT | 81.6 | 57.4 | 52.0 | 45.2 | 63.6 | 43.8 | 78.8 | 79.2 | 61.1 | 63.8 | 76.2 | 62.8 | 52.7 | 74.6 | 58.3 | 7.1 | 62.6 | |
1-shot | 92.5 | 66.7 | 62.5 | 62.5 | 59.4 | 58.3 | 70.8 | 82.5 | 66.7 | 77.5 | 70.8 | 78.8 | 37.5 | 50.0 | 75.0 | 37.5 | 71.6 | |
2-shot | 87.5 | 62.5 | 66.7 | 75.0 | 68.8 | 54.2 | 75.0 | 92.5 | 83.3 | 75.0 | 75.0 | 80.8 | 50.0 | 68.8 | 50.0 | 37.5 | 75.0 | |
Rule Hint | 82.5 | 62.5 | 62.5 | 50.0 | 78.1 | 29.2 | 70.8 | 87.5 | 66.7 | 60.0 | 70.8 | 76.9 | 25.0 | 93.8 | 50.0 | 12.5 | 68.5 |
Model | Prompt | ||||
Base | CoT | 1-shot | 2-shot | Rule Hint | |
Ernie-turbo | 4.9 | 2.4 | 6.7 | 7.9 | 5.5 |
Minimax | 31.3 | 20.7 | 40.3 | 44.3 | 35.7 |
Spark | 16.3 | 22.4 | 10.7 | 25.3 | 11.5 |
GPT4 | 65.0 | 59.3 | 59.9 | 64.1 | 57.2 |
GPT4-turbo | 54.2 | 50.7 | 49.4 | 53.8 | 49.4 |
Human | 95.7 | - | - | - | - |
5.1 Close Sourced LLMs’ Performance
In this section, we analyze the performance of closed-source LLMs in various tasks involving structure-rich text and multiple languages. Table 5 displays the overall performance on the complete StrucText-Eval dataset, while Table 2 compares both LLM and human performance specifically on the StrucText-Eval-Hard subset.
5.1.1 GPT4 Series Outperform All Other LLMs in Structure-Rich Text Understanding
The experiments revealed that GPT4 consistently outperformed other models across most tasks and languages. Specifically, GPT-4 achieved an average accuracy of 78.4 in all languages and tasks, significantly higher than the next best model, GPT4-turbo, which had an accuracy of 69.3. The marked disparity in performance between GPT4 Series and other LLMs underscores its robustness in tasks requiring deep semantic understanding and complex data manipulation. This demonstrates the potential of advanced models like GPT-4 in environments demanding high accuracy and versatility while also indicating the need for further development in other models to close the performance gap.
5.1.2 Current LLMs Cannot Truly Understand Structure-Rich Expressions
We conclude that although the performance of LLMs improves with more training data, they still fall short of truly comprehending structured language expressions, as evidenced by two observations. Firstly, evaluation results indicate that models from various companies exhibit distinct performance rankings across different languages. Notably, there is consistently lower performance in Tree structures, which are entirely composed of custom languages. This suggests that current LLMs are incapable of understanding complex, custom-structured expressions. Furthermore, results from utilizing Rule Hints generally underperform relative to the Base Prompt, implying that providing models with additional parsing rules may impair performance. In contrast, the 2-shot method consistently enhances the overall performance of LLMs, indicating that while LLMs are able to grasp individual examples, they do not fully understand the underlying structures or the rationale behind them.
5.1.3 Procedural Task is Watershed For LLMs
An in-depth analysis of the performance data from several LLMs demonstrates a clear demarcation in their ability to handle procedural tasks, which require navigating and manipulating structured data. Tasks emphasising procedural complexity, such as “PathCompose” and “Tree.Height”, present significant challenges to these models. For example, in the “PathCompose” task, GPT-4 shows a superior performance with an accuracy rate of 79.2% in its base setting, compared to Minimax, which achieves only 25.0%. Similarly, in the “Tree.Height” task, GPT-4 maintains a robust performance with 37.5% accuracy, whereas Minimax struggles, recording 0% in its base prompt setting.
5.1.4 Impact of Data Length and Complexity on LLM Performance
Fig. 5 illustrates how variations in Data Length and Complexity affect the performance of GPT4-Turbot. Specifically, we quantify Data Length and Complexity using the character length of References and the number of Nodes. It is evident that increasing the Node count while maintaining a Reference length under approximately 1000 tends to degrade performance, whereas increasing the Node count with a Reference length over 1000 enhances performance. This suggests that longer References complicate model comprehension, but additional structured expression aids understanding when Reference lengths are extensive. Conversely, when Reference lengths are shorter, introducing more structured expressions may introduce noise and reduce comprehension.
Fig. 5 displays the impact of different Reference lengths and Node counts on GPT4-Turbo across various languages and tasks. Significant impacts are noted in Node.Depth, PathCompose, and Statistics tasks, as well as in LaTeX, Markdown, and Tree languages, where increasing Reference length and Node count adversely affects performance. This could be due to the absence of complex structural data in these real-world applications, suggesting that training data did not contain excessive complexity, making additional nodes and References increase the complexity LLMs must manage, thereby affecting accuracy and efficiency. However, in the PathWalk task, while increasing the Node count degrades performance, a greater number of References exacerbates this effect by causing information overload and complicating path tracking. Conversely, increasing Reference length and Node count showed minimal impact on JSON, XML, and YAML, which is attributed to the frequent use and, thus, better generalization of GPT4-Turbo in these well-trained languages.
5.2 Performance Analysis of Finetuned Large Language Models
This section analyzes the enhancement of LLMs’ understanding of structure-rich text through finetuning. Fig. 7 and Fig. 7 illustrate the performance changes under different finetuning strategies. Notably, the initial accuracy of the Vanilla Llama-7B-v1 is 0%, which is reflected by the y-axis starting at 0% at 0 steps in all figures.
5.2.1 Finetuning Enhance LLM’s Performance in All Languages
As depicted in Fig. 7, finetuning improves the models’ comprehension of all structured languages. Although the base finetuning strategy may compromise some capabilities in general free-text language, it notably enhances the understanding of structured data, particularly in YAML and Tree languages, achieving performance comparable to GPT4-Turbo. This indirectly validates the rationality of the task settings for Tree language, as appropriate training enables LLMs to master this custom-designed language effectively. In contrast, while preserving capabilities in free-text language, the mix-fine-tuning strategy offers limited enhancement in the comprehension of structured data, with minimal improvements in Tree language. This may be due to the significant differences between the representation of Tree language and general text, making it challenging to enhance capabilities in both areas simultaneously with a 7B parameter model.
5.2.2 Not All Tasks can be Enhanced by FineTuning
Fig. 7 shows that changes in task performance are more pronounced than changes in language performance. Base finetuning provides the most substantial and consistent performance improvements across most tasks. As training progresses, model performance generally shows a continuous improvement trend. However, other finetuning methods do not correlate positively with increased training steps and performance enhancements in certain tasks, such as Join and Statistics. Additionally, tasks like Syntax quickly stabilize after a few training rounds, indicating that LLMs can rapidly adapt to and understand the grammatical rules of specific structured languages. This swift convergence may highlight the direct impact of finetuning strategies on the models’ comprehension of specific language structures.
@@ 5.3 Case Study
We illustrate the assessment setup and specific evaluation methods of StrucText-Eval through two exemplary cases in Fig. 8.
In the Text Retrieval task under the JSON language, the model is required to extract the most deeply nested objects from a complex nested structure and directly output the dictionary type in free-text format. The GPT4-Turbo model accurately identified the objects that met the criteria and matched the standard answer precisely, demonstrating its profound understanding of structured text. Although the Minimax model’s response was also judged as correct, it did not fully adhere to the preset output requirements. Given that many studies have explored the issue of compliance with output formats , our study focuses on assessing the LLM’s understanding of structured text, thereby considering Minimax’s response acceptable.
In the Join task under SQL language, the model is required to merge two tables and perform joint reasoning to output the number of records that meet specific criteria. The reference data was presented in a format that facilitates visualization. The GPT4-Turbo model, which without any further finetuning, failed to answer this question correctly. However, the model that underwent Base Finetuning exhibited progressive improvement: from initially incorrect responses at 300 training steps to gradually understanding the output format requirements at 1200 steps, and finally accurately solving the problem at 5100 steps. This progression illustrates the model’s incremental enhancement in comprehending complex SQL queries and database structures through appropriate finetuning.
@@ 6 Conclusion
We believe that the capability to directly interpret structural-rich text in a free-text format is an essential skill all LLMs require. In response, we have developed StrucText-Eval to evaluate this capability of LLMs. Our findings indicate that the proficiency of current LLMs in training on these structural-rich texts varies depending on user frequency, leading to markedly different outcomes when the same tasks are performed in various languages. Currently, LLMs’ understanding of structural-rich texts remains superficially tied to the training data, and these models lack a profound understanding of the structure itself. This deficiency becomes evident when LLMs encounter complex structures composed of common languages or need to parse structural-rich text by custom languages, resulting in significant performance degradation. Direct training on the dataset provided by StrucText-Eval effectively enhances the LLM’s understanding of structural-rich text. However, training solely on the StrucText-Eval dataset does not allow achieving GPT4-Turbo’s comprehensive capabilities on the test set, necessitating the collection of more extensive data and the optimization of training methodologies.
@@ 7 Limitation
This paper focuses on evaluating and enhancing the comprehension LLM’s abilities on structure-rich text by designing a benchmark called StrucText-Eval. However, StrucText-Eval includes only eight types of structured languages and encompasses a total of 29 different tasks. Given the vast array of actual structured languages and the myriad methodologies employed beyond these 29 types, StrucText-Eval can only partially represent the LLMs’ capacity to understand structure-rich text. Additionally, due to regional restrictions, we are unable to utilize some highly effective baseline LLMs, such as Gemini and Claude. Therefore, the conclusions drawn in this paper are based on the assumption that GPT-4 and GPT-4 Turbo represent the top-tier LLMs now.
@@ Appendix A Detail about Manual Works
Appendix B Statistics to StrucText-Eval
The detailed statistic information about StrucText-Eval is illustrated in Tab. 3.
Category | Task | Difficulty | #Samples | #Input | #Instruction | #Answer | #Node |
JSON | PathCompose | hard | 104 | 787.1 | 101.4 | 32.4 | 3 |
simple | 104 | 116.3 | 101.0 | 17.6 | 2 | ||
PathWalk | hard | 104 | 787.1 | 38.0 | 1.1 | 3 | |
simple | 104 | 116.3 | 38.0 | 1.0 | 2 | ||
Syntax | hard | 104 | 736.0 | 148.0 | 4.0 | 3 | |
simple | 104 | 109.0 | 148.0 | 4.0 | 2 | ||
TextRetrieval | hard | 124 | 786.6 | 105.1 | 156.4 | 3 | |
simple | 104 | 116.3 | 94.1 | 75.3 | 2 | ||
TextRetrieval_1∗ | hard | 104 | 787.1 | 161.0 | 315.4 | 3 | |
simple | 104 | 116.3 | 161.0 | 32.2 | 2 | ||
LaTeX | PathWalk | hard | 104 | 2726.1 | 179.1 | 484.0 | 2 |
simple | 104 | 277.2 | 160.3 | 145.1 | 1 | ||
TextRetrieval | hard | 104 | 2726.1 | 74.0 | 54.2 | 2 | |
simple | 104 | 277.2 | 74.0 | 16.1 | 1 | ||
TextRetrieval_1 | hard | 104 | 2726.1 | 62.0 | 57.1 | 2 | |
simple | 104 | 277.2 | 62.0 | 13.3 | 1 | ||
Markdown | PathWalk | hard | 104 | 2355.7 | 184.8 | 370.9 | 3 |
simple | 104 | 228.5 | 162.5 | 114.5 | 2 | ||
TextRetrieval | hard | 104 | 2355.7 | 73.0 | 52.5 | 3 | |
simple | 104 | 228.5 | 73.0 | 12.4 | 2 | ||
TextRetrieval_1 | hard | 104 | 2355.7 | 62.0 | 52.1 | 3 | |
simple | 104 | 228.5 | 62.0 | 12.2 | 2 | ||
ORG | PathWalk | hard | 104 | 2252.7 | 178.6 | 373.3 | 3 |
simple | 104 | 200.1 | 156.5 | 104.5 | 2 | ||
TextRetrieval | hard | 104 | 2252.7 | 73.0 | 52.0 | 3 | |
simple | 104 | 200.1 | 73.0 | 15.2 | 2 | ||
TextRetrieval_1 | hard | 104 | 2252.7 | 62.0 | 54.0 | 3 | |
simple | 104 | 200.1 | 62.0 | 12.3 | 2 | ||
SQL | SQL.Join | hard | 104 | 798.6 | 51.0 | 1.0 | 0 |
simple | 104 | 488.3 | 51.0 | 1.0 | 0 | ||
Statistics | hard | 104 | 798.6 | 49.9 | 1.3 | 0 | |
simple | 104 | 488.3 | 49.9 | 1.0 | 0 | ||
Statistics_1 | hard | 104 | 798.6 | 27.0 | 1.1 | 0 | |
simple | 104 | 488.3 | 27.0 | 1.0 | 0 | ||
TextRetrieval | hard | 104 | 798.6 | 43.8 | 4.3 | 0 | |
simple | 104 | 488.3 | 43.6 | 3.9 | 0 | ||
Tree | PathCompose | hard | 104 | 1745.7 | 84.0 | 25.0 | 8 |
simple | 104 | 555.6 | 83.8 | 11.3 | 4 | ||
Tree.Depth | hard | 104 | 1745.7 | 68.0 | 1.0 | 8 | |
simple | 104 | 555.6 | 67.9 | 1.0 | 4 | ||
Tree.Height | hard | 104 | 1745.7 | 156.0 | 1.0 | 8 | |
simple | 104 | 555.6 | 156.0 | 1.0 | 4 | ||
XML | Syntax | hard | 104 | 2135.7 | 147.0 | 4.0 | 5 |
simple | 104 | 2338.0 | 147.0 | 4.0 | 3 | ||
TextRetrieval | hard | 104 | 2621.6 | 54.5 | 17.7 | 5 | |
simple | 104 | 2910.4 | 52.5 | 11.9 | 3 | ||
TextRetrieval_1 | hard | 104 | 2621.6 | 81.9 | 110.4 | 5 | |
simple | 104 | 2910.4 | 83.8 | 132.9 | 3 | ||
YAML | PathCompose | hard | 104 | 631.7 | 101.4 | 32.4 | 6 |
simple | 104 | 82.2 | 101.0 | 19.6 | 4 | ||
PathWalk | hard | 104 | 631.7 | 38.0 | 1.0 | 6 | |
simple | 104 | 82.2 | 38.0 | 1.0 | 4 | ||
Syntax | hard | 104 | 725.8 | 148.0 | 4.0 | 6 | |
simple | 104 | 94.6 | 148.0 | 4.0 | 4 | ||
TextRetrieval | hard | 104 | 631.7 | 94.3 | 111.9 | 6 | |
simple | 104 | 82.2 | 94.0 | 51.9 | 4 | ||
TextRetrieval_1 | hard | 104 | 631.7 | 161.0 | 224.2 | 6 | |
simple | 104 | 82.2 | 161.0 | 22.1 | 4 |
Appendix C Detail Setting
All experiments and training process is carried out on a three 3090 GPUs service. The setting of API calling and finetuning is illustrated in Tab. 4 And the training loss is illustrated in Fig. 9(d).
Random Seed | ||||
42 | 42 | 42 | 42 | True |
AutoCausalLM | ||||
temperature | top_p | top_k | num_beams | max_new_token |
0.95 | 0.95 | 5 | 2 | 1 |
Training Detail | ||||
Batch Size | Max Length | All steps | Warm_up Steps | Learning Rate |
64 | 1024 | 5800 | 100 | 1e-5 |
Appendix D Detail Prompt

### v3

@@ 5 Analysis
Self-CoT | 52.2 | 40.6 | 49.2 | 39.0 | 66.5 | 43.2 | 36.6 | 55.2 | 40.9 | 40.2 | 39.7 | 53.0 | 77.3 | 65.9 | 52.3 | 36.4 | 48.5 | |
PS-CoT | 45.8 | 18.7 | 34.0 | 32.8 | 64.0 | 63.1 | 44.6 | 41.3 | 48.8 | 50.5 | 44.7 | 32.8 | 69.8 | 56.8 | 64.3 | 62.8 | 55.9 | |
[…]
Self-CoT | 93.6 | 71.4 | 69.7 | 54.8 | 96.0 | 84.1 | 87.1 | 95.9 | 88.6 | 67.9 | 86.4 | 85.7 | 97.7 | 93.2 | 77.3 | 97.7 | 76.7 | |
PS-CoT | 94.5 | 68.7 | 72.7 | 61.7 | 93.7 | 83.2 | 93.9 | 98.5 | 90.8 | 77.0 | 93.9 | 84.2 | 93.2 | 90.9 | 81.8 | 90.9 | 72.9 | |
[…]
Self-CoT | 87.7 | 62.2 | 74.2 | 62.2 | 95.5 | 75.8 | 78.5 | 90.8 | 87.9 | 73.2 | 63.6 | 83.4 | 100.0 | 90.9 | 59.1 | 88.6 | 67.1 | |
PS-CoT | 84.5 | 67.4 | 76.0 | 66.7 | 92.0 | 86.7 | 94.7 | 94.7 | 88.3 | 79.1 | 93.2 | 81.1 | 97.7 | 85.2 | 90.9 | 88.6 | 74.9 | |
[…]
Self-CoT | 56.5 | 35.1 | 40.3 | 36.6 | 34.7 | 15.9 | 33.7 | 54.3 | 28.8 | 49.2 | 64.4 | 43.9 | 6.8 | 23.9 | 13.6 | 13.6 | 8.1 | |
PS-CoT | 43.9 | 19.7 | 22.9 | 15.6 | 14.8 | 18.2 | 34.6 | 44.1 | 18.9 | 30.6 | 56.8 | 29.1 | 22.7 | 6.8 | 13.6 | 22.7 | 19.5 | |
[…]
Model | Prompt | |||
Base | w/ Hint | 3-Shot | Simple 3-Shot | |
GPT-4o-Turbo | 51.1 | 54.2 | 69.5 | 49.7 |
GPT-4o-Mini | 39.3 | 47.7 | 65.6 | 39.9 |
Gemini1.5-Pro | 11.2 | 15.7 | 53.0 | 12.5 |
Gemini1.5-Pro-Flash | 12.9 | 12.9 | 38.3 | 11.9 |
GLM-4-Plus | 47.3 | 50.9 | 65.8 | 51.7 |
GLM-4-Flash | 40.9 | 47.8 | 55.2 | 41.7 |
QWen-2-7B | 29.6 | 35.0 | 51.9 | 30.0 |
QWen-2-72B | 42.5 | 45.3 | 61.4 | 36.2 |
Llama-3.1-8B | 22.3 | 26.7 | 33.7 | 34.2 |
Llama-3.1-70B | 45.8 | 56.0 | 58.4 | 50.1 |
Llama-3.1-405B | 34.4 | 41.7 | 48.7 | 40.6 |
Mistral-0.2-7B | 7.0 | 9.5 | 21.0 | 6.9 |
Human | 92.6 | - | - | - |
5.1 Overall Performance in StrucText-Eval
The overall performance in StrucText-Eval is presented in Table 5, revealing significant variations in the performance of different models across various languages and tasks. For instance, the Qwen2-72B-Instruct model demonstrates optimal performance on JSON-formatted tasks with an 85.8% accuracy under the “Naive” prompt. It also achieves notable results in YAML and CSV tasks, with accuracies of 82.7% and 86.4%, respectively. In contrast, the Meta-Llama-3.1-8B-Instruct-Turbo model performs poorly under the same settings, achieving only 64.6% accuracy on LaTeX tasks. Manually injected hints (w/ hint) generally improve model performance, particularly in tasks requiring deep reasoning, such as those involving YAML and JSON. For example, the Meta-Llama-3.1-70B-Instruct-Turbo model’s accuracy improves from 75.4% under the “Naive” prompt to 84.9% with the “w/ Hint” strategy. However, with “Self-CoT” and “PS-CoT” prompts, specific models like Qwen2-7B-Instruct exhibit lower accuracy across multiple tasks, especially when handling complex structures such as XML and Tree data, performing significantly worse compared to other prompting methods.
These performance disparities can be primarily attributed to training sample biases and the influence of different prompting strategies. JSON, being a widely used format in internet data, is frequently encountered by many large models during training, leading to a pronounced advantage in handling JSON-formatted tasks—a clear manifestation of training sample bias. Moreover, the choice of prompting strategy directly affects a model’s inference capabilities. The “w/ Hint” method, which introduces human reasoning rules, compensates for the model’s limitations in reasoning through complex structures. Conversely, while the “Self-CoT” and “PS-CoT” approaches encourage step-by-step reasoning, they often result in logical inconsistencies and reasoning errors in complex tasks due to the requirement for autonomous generation of reasoning paths.
5.2 Overall Performance on StrucText-Eval Hard
Table 3 presents the performance of various models on the StrucText-Eval Hard dataset, characterized by more complex tasks with longer sequences and deeper structures. This complexity results in a significant performance decline across all models. For instance, the accuracy of the Qwen2-72B-Instruct model decreases from 78.4% to 65.0%, while the Meta-Llama-3.1-70B-Instruct-Turbo model’s accuracy drops sharply from 75.4% to 43.2%. Unlike the standard dataset, the Hard dataset demands more advanced reasoning skills, and even with the “w/ Hint” strategy, models achieve only limited improvements, in contrast to the substantial gains observed in more straightforward contexts. Notably, human accuracy on StrucText-Eval-Hard reaches 95.7%, significantly surpassing that of the best-performing large language models (LLMs), highlighting a considerable gap in models’ capabilities for structured reasoning.
This performance gap can be primarily attributed to biases in training data and the limitations of current prompting methods. The StrucText-Eval Hard dataset, with increased question complexity and depth, requires models to possess enhanced abstraction abilities and a deeper understanding of complex structures. However, most models are trained on relatively more straightforward structured text, which makes them less effective when tackling deeply nested reasoning tasks. Additionally, prompting methods like “w/ Hint” fail to achieve human-level understanding in multi-layered scenarios. The differences in prompting methods become more pronounced with increased complexity; more straightforward methods, such as Self-CoT, need to be revised for guiding models through multi-step reasoning in these challenging contexts. While the “3-shot demonstration” approach significantly improves model performance, the simpler “simple 3-shot” method, despite following similar reasoning rules, fails to match the former due to its insufficient complexity.
5.3 Few-Shot Demonstration on Structural Text Inference
Figure 4 demonstrates that model performance improves with an increasing number of demonstrations under Few-Shot settings. In the 3-shot scenario, GPT-4 achieves an accuracy of 69.5%, significantly outperforming models like Gemini-Pro-Flash and Mistral, which remain around 21% or lower. The Qwen-2-72B-Instruct model shows steady improvement as more examples are provided, although it continues to trail behind GPT-4. Generally, performance increases from 1-shot to 3-shot, but the gains become less pronounced at 5-shot, with some models showing overfitting. In contrast, the performance of CoT and PS approaches remains less consistent as the number of demonstrations increases.
This trend suggests that a more significant number of examples helps models to understand problem structures and reasoning processes better, thereby enhancing their inference capabilities. However, providing too many examples can lead to models overfitting to specific patterns, which diminishes their ability to generalize to new tasks. The quality and diversity of examples are critical—high-quality examples can guide practical reasoning, while poor examples may mislead the models. While few-shot learning enhances model adaptability, those with limited pretraining data or lower parameter counts may struggle to capitalize on this approach entirely. For CoT and PS methods, the reasoning process requires additional steps, which means that simply increasing the number of few-shot demonstrations does not consistently yield performance improvements.
5.4 Model Performance Across Different Difficulty Levels, Languages, and Tasks
Figure 5 illustrates the performance variations of models across different languages and tasks. The two figures on the left reveal that, while numerical differences exist among models, including GPT models, they exhibit a consistent trend: Increasing the reference’s depth and width results in a significant decline in performance. Notably, all models show a high variance in performance when the depth and width are limited, suggesting that the StrucText-Eval Test suite effectively distinguishes the capabilities of most models under these conditions.
However, for GPT models, substantial variance in performance is observed only when the depth and width increase significantly, indicating that the StrucText-Eval-Hard Test suite is necessary to better differentiate the performance of more advanced models. Additionally, there is considerable variance in model performance across different languages and tasks, suggesting substantial differences in models’ proficiency in handling various linguistic and task-specific challenges. This discrepancy is likely due to biases in training samples and the varying difficulty levels of those samples, as suggested by earlier analyses.
5.5 Correlation Between Different Metrics
Figure 6 presents the correlations between various evaluation metrics. The high correlation between Human Judge and GPT-4o Judge (0.9937) indicates a strong alignment between GPT-4o’s automated assessments and human evaluations. Although Exact Match exhibits a notable correlation with Human Judge (0.9501), its stringent criteria often result in scores significantly lower than those of human evaluators, making it less suitable for capturing the diversity and naturalness of model outputs. Among the metrics, RougeL stands out with a correlation of 0.9932 with Human Judge, demonstrating its effectiveness in capturing surface-level textual similarity while maintaining high consistency with human judgments. Compared to the more rigid Exact Match and the relatively lower correlation of BLEU, RougeL offers a better balance between textual similarity and evaluation accuracy.
@@ 5.6 Case Study
Two case studies illustrate the evaluation setup of StrucText-Eval (Figure 7). In the JSON-based Text Retrieval task, GPT4-Turbo accurately identified deeply nested objects and adhered to the free-text format for outputting dictionary types, reflecting its firm grasp of structured text. Minimax also produced a correct answer but deviated from the prescribed format, a common issue explored in existing research. In contrast, GPT4-Turbo initially failed to merge two tables and deduce the correct record count without fine-tuning in the SQL-based Join task. However, a finetuned model steadily improved, achieving the correct solution after 5100 training steps. This progression demonstrates the importance of task-specific fine-tuning in enhancing models’ capabilities in handling complex SQL queries and database structures.
@@ 6 Conclusion
The capability to directly interpret structural-rich text in a free-text format is an essential skill all LLMs require. In response, we have developed StrucText-Eval to evaluate this capability of LLMs. Our findings indicate that the proficiency of current LLMs in training on these structural-rich texts varies depending on user frequency, leading to markedly different outcomes when the same tasks are performed in various languages. LLMs’ understanding of structural-rich texts remains superficially tied to the training data, and these models need a profound understanding of the structure itself. This deficiency becomes evident when LLMs encounter complex structures composed of common languages or need to parse structural-rich text by custom languages, resulting in significant performance degradation.
@@ 7 Limitation
This paper focuses on evaluating LLM’s reasoning abilities on structure-rich text by designing a benchmark called StrucText-Eval. However, StrucText-Eval includes only eight types of structured languages and encompasses a total of 29 different tasks. Given the vast array of actual structured languages and the myriad methodologies employed beyond these 29 types, StrucText-Eval can only partially represent the LLMs’ capacity to understand structure-rich text. Additionally, due to regional restrictions, we are unable to utilize some highly effective baseline LLMs, such as Gemini and Claude. Therefore, the conclusions drawn in this paper are based on the assumption that GPT-4 and GPT-4 Turbo represent the top-tier LLMs now.
@@ Appendix A Detail about Manual Works
Appendix B Other Metrics
Given the substantial expense in evaluating all results using multiple metrics, we selected a subset of 300 test results for each model on the StrucText-Hard dataset, using a naive prompting method for assessment. The complete evaluation results are presented in Table 4.
Model | Human | GPT-4o | 4o-Mini | RougeL | BLEU | EM |
GPT-4o-Turbo | 56.13 | 55.75 | 51.00 | 51.1 | 45.94 | 40.31 |
GPT-4o-Mini | 36.15 | 36.02 | 40.73 | 39.3 | 46.08 | 33.93 |
Gemini1.5-Pro | 12.39 | 12.80 | 10.62 | 11.2 | 12.60 | 8.75 |
Gemini1.5-Pro-Flash | 13.83 | 13.19 | 12.96 | 12.9 | 14.01 | 9.67 |
GLM-4-Plus | 52.90 | 52.62 | 46.02 | 47.3 | 32.75 | 38.27 |
GLM-4-Flash | 41.50 | 41.34 | 38.99 | 40.9 | 37.43 | 34.80 |
QWen-2-7B | 32.95 | 31.99 | 30.10 | 29.6 | 27.98 | 18.70 |
QWen-2-72B | 40.87 | 38.66 | 31.24 | 42.5 | 37.76 | 35.67 |
Llama-3.1-8B | 21.78 | 21.98 | 22.36 | 22.3 | 20.88 | 14.75 |
Llama-3.1-70B | 46.64 | 41.38 | 40.83 | 45.8 | 41.50 | 27.46 |
Llama-3.1-405B | 35.01 | 35.97 | 35.88 | 34.4 | 28.00 | 21.29 |
Mistral-0.2-7B | 7.85 | 7.33 | 7.32 | 7.0 | 5.09 | 4.47 |
Appendix C Detail Prompt